• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name bogdanw

    Log entry time 08:06:51 on April 4,2010

    Entry number 316110

    keyword=shift summary



    Shift Summary - Owl shift







    Shift Summary    Owl Shift    April 4 2010    (PREx)




    Bogdan Wojtsekhowski (SL)

    Jon Wexler (TO+3rd)




    (one needs character uncoding in Browser set to Unicode to view properly)



    00:00 Status at start of shift:


    Tune Y counting mode 0.5 uA on superthin target



    Tune X for Left arm

    Tube B for both arms

    01:45 Control access to put AT detector to integrating mode (Parity DAQ)



    02:10 Control access completed. Move the target to EMPTY



    02:10 Both arms has HV for VDC OFF, S0 OFF, Upper and Lower scint OFF , Lumi is ON



    02:30 Beam is 5 uA, positions OK, mode 240 Hz, reporting delay 16


    02:45 Beam is 5 uA, raster size is 4x4 however position is shifted up 3 mm, move to BeO


    02:55 Beam is 10 uA, some light is observed when laser ON/OFF


    03:05 Target BeO, Beam is 10 nA = att. 65 slit 17


    03:15 Left arm HV upper scin. T5, rate is 25,000 Hz


    Investigation why the rate is high, Rate vs raster current is OK.
    Use att. 55 so rate is 1.5 kHz
    03:45 Move to Lead#3 target, asked 3 uA


    A_T scan

    05:50 Q1 trip, He level recovery take long



    06:50 Q1 recover, John continue A_T scan




    Runs Taken



    Pedestals


    ---